the velocity of a particle moving along x axis is given by -


v = 32 t ^2 - 2t^3 m/s. what is the accelaration of the particle after maximum displacemen is achieved?

Grade:12

2 Answers

Chetan Mandayam Nayakar
12 years ago

at maximum displacement,dx/dt=v=0, 32t^2=2t^3,t=16s, a=dv/dt=64t-6t2=64*16-6*162

 

goutham 1995 gumm
12 years ago

in this case, if we compute the answer is  -512 m/sec^2. But the actual answer is -64m/sec^2. can someone tell how?